# Neuromorphic Control of a Pendulum

### v1

@@ VII Phase Control
Figure 10: Phase response curve of an HCO. The curve indicates the phase shift resulting from a fixed pulse ($w=0.05$ s and $P=0.3$) as a function of the timing of the pulse along the limit cycle.
Fig. 11 illustrates the stabilizing property of a proportional phase control on the behavior illustrated in fig. 7 (right plot). The phase control stabilizes the large oscillations to a constant period. We selected $q_{\rm{phase}}=-1$ rad such that small oscillations are prevented. To prevent braking during the transient, a sensory event is ignored if the pendulum’s velocity is negative.
Figure 11: The simulation of fig. 7 (right plot) with the addition of phase control. Large oscillations are now entrained. There is a slight change in frequency due to a constant steady-state error in phase. The current pulses to the A and B neurons are shown in green and purple respectively, and are scaled for clarity. The two plots are from the same simulation; they show both the initial part of the transient and the steady-state behaviour.
VIII Conclusion
We have presented a neuromorphic, event-based framework for controlling the oscillations of a mechanical pendulum. A simple neuronal network is used to generate rhythmic events capable of entraining the rhythmic behaviors of the pendulum. Adaptive control regulates the entrainment. Phase control enlarges the basin of attraction of a desired oscillation.
The event-based control strategy is inherently robust to model and controller uncertainties because of the high impedance but highly localised-in-time nature of the interactions.
The approach of this letter is general, and applies to a wide range of mechanical systems. Indeed, one need not stray too far from the pendulum itself to find familiar robotic or biomechanical rhythms: the ’inverted pendulum’ model lies at the foundation of the literature on walking [31, 32]. It consists of two rigid legs, one a pendulum, the other an inverted pendulum. To include a second state, running, add springs [33]. Meanwhile, multiple HCOs can be chained together to drive an amphibious lamprey-like robot that walks and swims [4]. Compliance is another topic of interest; owing to the high localised but low average impedance, pulse-based actuation suggests a path to overcoming the stiffness-compliance trade-off [34]. We hope that this letter will motivate such investigations in more challenging robotic designs.

### v3

@@ 6 Phase Control
Figure 8: Left: A small inhibitory pulse (green) causes a phase advance or delay of the subsequent burst. The blue curve shows $v_{A}$ for an isolated HCO. The blue dashed curve shows the behaviour of $v_{A}$ in the absence of the pulse. Right: Phase response curve of an HCO. The curve indicates the phase shift resulting from a fixed pulse ($w=0.05$ and $P=0.3$) as a function of the timing of the pulse along the limit cycle.
Fig. 9 illustrates the stabilizing property of a proportional phase control. The initial conditions are those of Fig. 7 (right plot), but the phase control forces convergence to the high- rather than low-energy state. The steady-state frequency is slightly different to that obtained with feedforward entrainment, and varies with $q_{p}$.
Figure 9: The simulation of Fig. 7 (right plot) with the addition of phase control, forcing convergence to the high- rather than the low-energy state. Left: the initial response. The current pulses to the A neurons are shown in purple, and are scaled for clarity. Each pink marker indicates the start of an actuation event. Right: the energy transferred per event, $E_{i}$, showing convergence to a steady-state oscillation. The pink markers match those of the left plot.
7 Neuromorphic Adaptive Control
Online adaptation of the control parameters can be used as well as phase control in order to regulate an oscillation.
Robust adaptive regulation of the control parameters has been previously studied in [12, 13, 32, 33]. This work relies on the property that the proposed neuronal architecture inherently satisfies the standard assumptions of adaptive control: the system is minimum phase, relative degree one, and linear in the parameters. Adaptive control is analogous to the biological concept of neuromodulation [34].
We consider the online adaptation of the two parameters ${g_{us}^{+}}$ and ${g_{s}^{-}}$, which regulate the neural frequency and burst size respectively. We use these to regulate the frequency and amplitude of small oscillations in ANTI-PHASE. The online adaptation is implemented via the integral action of a prediction error computed from the sensory events.
For a reference frequency ${\omega_{\rm{ref}}}$, we predict that a zero crossing $q(t)=0$ will occur every $\pi/{\omega_{\rm{ref}}}$ seconds. A correction $p_{\omega}$ is added to the parameter ${g_{us}^{+}}$ at each zero crossing, with a sign determined by the sign of the prediction error and an amplitude proportional to its magnitude.
Likewise, for a reference amplitude $A_{ref}$ for the pendulum’s oscillation, we predict that an event $q={A_{\rm{ref}}}$ occurs $\pi/2{\omega_{\rm{ref}}}$ seconds after each $q=0$ event. If no such event occurs before the next zero crossing, we apply a fixed correction $p_{A}$ to the parameter ${g_{s}^{-}}$. Otherwise, the correction is applied at the event $q={A_{\rm{ref}}}$ with a sign and amplitude determined by the prediction error.
The simple adaption law described above can be regarded as a phase integral control. In practice, the parameter variation is confined to a finite range such that bursting is preserved.
Fig. 10 illustrates parameter convergence for the reference values illustrated in Figure 6.
Figure 10: The gains tuned by the adaptive controller, when used to obtain the parameters in Fig. 6. The legend gives the values of ${A_{\rm{ref}}}$ in radians. In each case, ${\omega_{\rm{ref}}}$ is the fixed frequency chosen in Section 5.
8 Discussion
We have presented a neuromorphic, event-based framework for controlling the oscillations of a mechanical pendulum. A simple neuromorphic controller was used to generate actuator events capable of entraining the pendulum oscillation while regulating its energy. The feedforward controller exhibits bistability between low-energy and high-energy oscillations in the underdamped regime. A proportional phase control was designed to stabilise the high energy oscillation only. Finally, an integral phase control was designed to regulate oscillations at a given frequency and amplitude. While no formal convergence analysis was included in the letter, the design was supported by existing analysis and system properties that enable regulation by simple feedback controllers (minimum phase and relative degree one).
The event-based nature of the designed controller offers a number of potential advantages. Prime and foremost, the energy exchange between the systems is confined to the events. As a consequence, the temporal sparsity of the events is a direct measure of the energy efficiency of the design. High impedance control during the events is an inherent source of robustness to model uncertainty. Temporal sparsity of the events simultaneously ensures low impedance control when averaged over time. In this way, event-based control could potentially combine the benefits of soft actuation with the high impedance requirements of robust control, suggesting a path to overcoming the classical stiffness-compliance trade-off [35]. The considered two-motor architecture also suggests how actuation events can be easily distributed over possibly many actuators, with phase relationships ensured by the architecture of the controller.
The purpose of the letter was to illustrate the potential of neuromorphic control in a simple mechanical system. We hope that it will stimulate further theoretical and practical research in the control of robotic or biomechanical rhythms. It would seems natural to explore the potential of neuromorphic control architectures in robotic designs that have already exploited rhythmicity, for instance [4].